OpenStudy (anonymous):

How do I solve this question 2/3y+7=17?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {2 \over 3y} + 7 = 17 }\) Is this the question?

OpenStudy (anonymous):

No the Y is not with the fraction

Parth (parthkohli):

Ok.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {2 \over 3}y + 7 = 17 }\) \( \color{Black}{\Rightarrow \Large {2 \over 3}y = 10 }\) \( \color{Black}{\Rightarrow y = {10} \times \Large {3 \over 2} }\) \( \color{Black}{\Rightarrow y = 15 }\)

OpenStudy (anonymous):

How do you get 15 from 10 x 3/2?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 10 \times \Large{3 \over 2} }\) \( \color{Black}{\Rightarrow \Large{10 \times 3 \over 2} }\) \( \color{Black}{\Rightarrow \Large{30 \over 2} }\) \( \color{Black}{\Rightarrow 15 }\)

OpenStudy (anonymous):

I see. Thank You!

Parth (parthkohli):

:) pleasure

OpenStudy (anonymous):

wait so why did you switch the 2/3 around?

OpenStudy (eyust707):



OpenStudy (eyust707):